OpenStudy (anonymous):

Soo will rent a car for the weekend. She can choose one of two plans. The first plan has an initial fee of$44and costs an additional$0.14per mile driven. The second plan has an initial fee of$51and costs an additional$0.12per mile driven. for what amount of driving would the 2 plans cost the same? how would i calculate that and what is the cost when the 2 plans cost the same

OpenStudy (anonymous):

Help someone how do I do this?

OpenStudy (anonymous):

@surjithayer would 168 be the price for both

OpenStudy (anonymous):

12x14=168 and 14x12=168

OpenStudy (anonymous):

or would the cots be 1.68

OpenStudy (anonymous):

im lost please someone

OpenStudy (anonymous):

@One098 Can you help me please I'm lost

OpenStudy (anonymous):

?

OpenStudy (texaschic101):

m stands for # of miles 44 + 0.14m = 51 + 0.12m --- set the 2 plans equal to each other 0.14m - 0.12m = 51 - 44 -- get the m terms on one side 0.02m = 7 m = 7/0.02 m = 350 The 2 plans will cost the same at 350 miles check.. 44 + 0.14(350) = 51 + 0.12(350) 44 + 49 = 51 + 42 93 = 93 (correct) And when driving 350 miles, each plan costs $ 93

OpenStudy (anonymous):

Thanks @texaschic101 a lil fast but i think i get it

OpenStudy (anonymous):

one last question if u could help @texaschic101 To rent a certain meeting room, a college charges a reservation fee of $48 and an additional fee of $9.40 per hour. The math club wants to spend less than $113.80 on renting the meeting room. What are the possible amounts of time for which they could rent the meeting room? Use t for the number of hours the meeting room is rented, and solve your inequality for t .

OpenStudy (texaschic101):

48 + 9.40t < 113.80 9.40t < 113.80 - 48 9.40t < 65.80 t < 65.80/9.40 t < 7 They can rent the room for anything UNDER 7 hrs

OpenStudy (anonymous):

@texaschic101 can you explain that one more thoroughly

OpenStudy (anonymous):

Let the distance driven be x miles in two plans. 44+0.14x=51+0.12x 0.14x-0.12x=51-44 0.02 x=7 \[x=\frac{ 7 }{ 0.02 }=7\times \frac{ 100 }{ 2 }=7\times50=350 miles\]

OpenStudy (anonymous):

cost in each case\[=44+0.14 x=44+\frac{ 14 }{ 100 }\times350=44+49=393 $\]

OpenStudy (anonymous):

correction 44+49=93 $